### v1

# Bayesian Despeckling of Structured Sources

Abstract
Speckle noise is a fundamental challenge in coherent imaging systems, significantly degrading image quality. Over the past decades, numerous despeckling algorithms have been developed for applications such as Synthetic Aperture Radar (SAR) and digital holography. In this paper, we aim to establish a theoretically grounded approach to despeckling. We propose a method applicable to general structured stationary stochastic sources. We demonstrate the effectiveness of the proposed method on piecewise constant sources. Additionally, we theoretically derive a lower bound on the despeckling performance for such sources. The proposed depseckler applied to the 1-Markov structured sources achieves better reconstruction performance with no strong simplification of the ground truth signal model or speckle noise.
1 Introduction
1.1 Problem statement
Multiplicative noise, commonly referred to as speckle noise, poses a significant challenge in coherent imaging systems such as synthetic aperture radar [1], optical coherence tomography [2], and digital holography [3]. The inherent non-linearity of the multiplicative noise model complicates the analysis and design of optimal despeckling algorithms—denoising methods tailored to address multiplicative noise and recover the underlying signal—even for relatively simple structured sources [4].
Mathematically, the despeckling problem can be formulated as follows. Consider a stationary stochastic process ${\bf X}=\{X_{i}\}_{i\in\mathbb{N}}$, $X_{i}\in\mathcal{X}$, observed under a multiplicative noise model, $Y_{i}=X_{i}W_{i}$, where $W_{i}$ represents the speckle noise. In most coherent imaging applications, the speckle noise process is assumed to be fully developed and is therefore modeled as Gaussian [5]. Here, we assume that $\{W_{i}\}_{i}$ are independent and identically distributed as standard normal random variables, $\mathcal{N}(0,1)$. The despeckling (or denoising under speckle noise model) goal is to recover $X^{n}$ from speckle-corrupted measurements $Y^{n}$.
In a Bayesian framework, given the source distribution $p(x^{n})$, the optimal MMSE despeckler lets ${\hat{X}}^{n}=\operatorname{E}[X^{n}|Y^{n}]$. However, even for additive noise, direct computation of $\operatorname{E}[X^{n}|Y^{n}]$ is often intractable and highly challenging. Additionally, the source distribution $p(x^{n})$ is rarely accessible in practice, and typically, only samples from this distribution are available. These challenges raise a fundamental question:
Question. Can we design a theoretically founded, computationally efficient framework for despeckling that applies to a broad class of structured sources?
1.2 Related work
Theoretically derived Bayesian despeckling methods can be broadly categorized into two approaches: adaptive linear minimum mean squared error (MMSE) filtering and maximum a posteriori (MAP) estimation [4, 1]. In early work, [6] proposed an affine approximation to the nonlinear speckle model, minimizing the $\ell_{2}$ norm and matching first-order moments to derive the MMSE estimator. A fully driven linear MMSE approach was later introduced in [7], avoiding the approximation errors of the affine model. Both methods rely on local statistics calculated over pre-defined windows, yielding comparable despeckling performance in practice [8]. A refinement for images was proposed in [9], leveraging the local gradient to adaptively redefine neighborhoods for improved statistical estimation.
Beyond linear MMSE methods, MAP estimators have been extensively studied, often requiring strong assumptions about the signal prior distribution. For example, [10] considered a MAP framework with a gamma prior, in contrast to the Gaussian prior used in [11]. These methods generally assume local stationarity of the signal and parameterize its distribution using moments computed within a local window.
Other approaches explore regularization-based techniques, such as total variation (TV) minimization [12]. For speckle noise, [13] combines a gamma likelihood fidelity term with the TV regularizer, though the inherent nonconvexity of the speckle model limits its practical utility. A log-transform-based linearization of the speckle model was proposed in [14], where the additive noise is denoised and rescaled back by exponentiation. However, this approach suffers from degraded performance due to mismatches between the log and signal domains.
Additionally, drawing inspiration from denoising methods designed for additive noise, various heuristic despeckling approaches—such as non-local means [15], SAR-BM3D [16], and neural network-based solutions [17]—have been widely adopted, particularly for high-resolution imaging tasks. For instance, the DnCNN architecture [18] has been trained for despeckling images, both with the log-transformation [19] and without it [20]. Also, self-supervised despeckling algorithms, based on the Noise2Noise [21] and Noise2Void [22] frameworks originally developed for additive denoising, are extended to the despeckling problem in [23] and [24], respectively. State-of-the-art generative models have also been explored for addressing the image despeckling problem. These include the use of generative adversarial networks as demonstrated in [25] and diffusion-based probabilistic models as explored in [26].
1.3 Our contribution
We propose Bayesian Despeckling via QMAP (BD-QMAP), a novel despeckling algorithm inspired by the Quantized Maximum a Posteriori (Q-MAP) estimator [27], originally developed for additive noise. QMAP defines the estimator as the minimizer of a cost function over the space of all possible solutions within the signal’s support space $\mathcal{X}^{n}$. The cost function consists of two terms: a negative log-likelihood term capturing the noise model and a regularization term enforcing distributional similarity between the quantized reconstruction and quantized ground-truth data. By leveraging learned statistics, QMAP assigns weights to unique realizations of the reconstruction, effectively reducing the impact of irrelevant solutions.
Building on these ideas, BD-QMAP adapts QMAP principles to multiplicative noise, tailoring the framework for general structured sources. To clarify its operation, we simplify the formulation for classic structured sources, including memoryless sources with a mixture of continuous and discrete components and piecewise-constant sources. We establish a theoretical lower bound on BD-QMAP’s performance for piecewise-constant first-order Markov processes. Experimental results demonstrate that BD-QMAP achieves state-of-the-art performance on piecewise-constant sources modeled as first-order Markov processes.
1.4 Notations and definitions
Finite sets are denoted by calligraphic letters. For a finite set $\mathcal{A}$, $|\mathcal{A}|$ denotes its size. For $b\in\mathbb{N}^{+}$, the $b$-bit quantized version of $x\in\mathbb{R}$ is denoted as $[x]_{b}$, defined as $[x]_{b}=2^{-b}\lfloor 2^{b}x\rfloor$. For $x^{k}\in\mathcal{X}^{k}$, $[x^{k}]_{b}$ denotes the element-wise $b$-bit quantization of $x^{k}$. The $b$-bit quantized version of $\mathcal{X}\subset\mathbb{R}$ is denoted by $\mathcal{X}_{b}$, defined as
$\mathcal{X}_{b}=\{[x]_{b}:\;x\in\mathcal{X}\}.$
Consider sequence $u^{n}\in\mathcal{U}^{n}$, with finite alphabet $\mathcal{U}$ ($|\mathcal{U}|<\infty$). The $k$-th order empirical distribution of $u^{n}$, $\hat{p}^{k}(\cdot|u^{n})$, is defined as follows. For $a^{k}\in\mathcal{U}^{k}$,
$\hat{p}^{k}(a^{k}|u^{n})=\frac{1}{n-k+1}\sum_{i=1}^{n-k+1}\mathds{1}_{u_{i}^{i+k-1}=a^{k}}.$ (1)
1.5 Paper organization
This paper is organized as follows: Section 2 introduces the BD-QMAP despeckler. Section 3 discusses its application to two classical structured sources and derives a lower bound for the MSE of the piecewise constant source. Section 5 describes numerical experiments, including the algorithm’s implementation and a performance comparison with other despeckling methods.
2 Bayesian Despeckling via QMAP
Consider the despeckling problem, where the goal is to recover $X^{n}$ from noisy measurements $Y^{n}=X^{n}W^{n}$. Here, $X^{n}$ follows a known prior distribution $p(x^{n})$ and $W^{n}$ is i.i.d. $\mathcal{N}(0,\sigma_{w}^{2})$. In this formulation, we neglect the effect of additive noise, focusing solely on the multiplicative noise model, as is commonly done in coherent imaging applications. Note that under this model, the variance of speckle noise does not play a role, as one can always scale the measurements by $1/\sigma_{w}$. Therefore, in the remainder of the paper, without loss of generality, we assume that $\sigma_{w}^{2}=1$.
Inspired by the Q-MAP denoiser, introduced for additive noise [27], we propose BD-QMAP, a novel despeckling algorithm that is applicable to general structured sources. To define this approach, first we review how the structure of the source is accounted for in the original Q-MAP algorithm. For $a^{k}\in\mathcal{U}_{b}^{k}$, define $w_{a^{k}}>0$ as
$w_{a^{k}}=-\log\operatorname{P}([X^{k}]_{b}=a^{k}),$
where the probability $P$ is computed with respect to the known distribution $p(x^{n})$. Here, $k\in\mathbb{N}^{+}$ and $b\in\mathbb{N}^{+}$ denote the memory parameter and the quantization level, respectively. Then, given weights $\mathbf{w}=(w_{a^{k}}:a^{k}\in\mathcal{U}_{b}^{k})$, the weight assigned to sequence $u^{n}\in\mathcal{U}^{n}$ is defined as
$c_{\mathbf{w}}(u^{n})=\sum_{a^{k}\in\mathcal{X}^{k}_{b}}w_{a^{k}}\hat{p}^{k}(a^{k}|u^{n})(a^{k}|[X^{k}]_{b}).$ (2)
Finally, the BD-QMAP method, recovers $X^{n}$ as $\hat{X}^{n,(k,b)}$ defined as
$\hat{X}^{n,(k,b)}=\operatorname*{arg\,min}_{u^{n}\in\mathcal{X}^{n}}{1\over n}\sum_{i=1}^{n}\big{(}\log u_{i}^{2}+\frac{Y_{i}^{2}}{u_{i}^{2}}\big{)}+\frac{\lambda}{b}c_{\mathbf{w}}(u^{n}).$ (3)
The cost function in (3) comprises of two terms. The first term represents a fidelity criterion derived from the negative likelihood of the observations given the signal under the multiplicative noise model. Note that given $X^{n}=x^{n}$, $Y^{n}\sim\mathcal{N}({\bf 0},X^{2})$, where $X=\mathop{\rm Diag}\nolimits({\bf x})$. Therefore,
$-\log p(y^{n}|x^{n})={1\over 2}\sum_{i=1}^{n}(\log x_{i}^{2}+{y_{i}^{2}\over x_{i}^{2}})+C.$
where $C$ is a constant not depending on $x^{n}$ or $y^{n}$. The second term in (3) imposes a prior on the ground-truth signal using the set of weights assigned to the quantized representation of the candidate reconstruction. The weights defined in (2) are a function of the signal’s known distribution. In other words, the weights are expected to summarize the source’s $n$-dimensional distribution characterized by $p(x^{n})$ into a finite number $|\mathcal{X}_{b}|^{k}$ of positive weights.
A key advantage of this approach is that it leads to a tractable approach for modeling and utilizing the sources structure. It can be shown that for structured sources, i.e., the sources with information dimension strictly less than one [28, 29], identified by having singularities in their distributions, e.g., spike and slap distribution, the weights ${\bf w}$ can be divided into two groups: 1) A small set of weights that identify the key structure of the source, and 2) a large set of weights that have negligible impact in the optimization [27].
Remark 1.
Solving the BD-QMAP optimization is challenging due to the nature of the cost function, which combines a non-convex loss term derived from the log-likelihood and a regularization term defined on the discretized space of sequences. To address this complexity, the optimization can be simplified by restricting the search space to $\mathcal{X}_{b}^{n}$, the space of quantized sequences. As explained in Section 5.1, this restriction enables the application of the Viterbi algorithm for efficient optimization. While this approach is suboptimal compared to solving the original problem, it offers significant computational advantages. Moreover, for cases such as the piecewise-constant source studied later, this simplification facilitates learning the structure of the jumps, allowing the original optimization to be solved in the continuous space with improved computational efficiency.
To gain deeper insight into the BD-QMAP optimization and the roles of its two terms, we examine two classic structured processes in the next section: structured i.i.d. sources and piecewise-constant sources. For each source, we derive a simplified form of the BD-QMAP despeckler optimization. Moreover, for the piecewise-constant source, we theoretically establish a lower bound on the expected mean squared error (MSE) achievable by the BD-QMAP method.
3 Analysis of BD-QMAP for classic structured sources
In the following, we focus on two classic structured source models and study BD-QMAP under each model.
3.1 Structured memoryless source
Consider an i.i.d process ${\bf X}$ such that $X_{i}\sim(1-q_{0})\delta_{x_{m}}+q_{0}{\rm Uniform}(x_{m},x_{M})$, Setting $k=1$ and using the results of [27, Sec. 3.1] to simplfy the second term in (3), the BD-QMAP algorithm can be written as
${\hat{X}^{n,(1,b)}=\arg\min_{u^{n}\in\mathcal{U}^{n}}\Big{[}{1\over n}\sum_{i=1}^{n}\big{(}\log u_{i}^{2}+\frac{Y_{i}^{2}}{u_{i}^{2}}\big{)}+\lambda(1+\gamma)(1-\hat{p}^{1}(x_{m}|[u^{n}]_{b}))\Big{]}},$
where $\gamma={1\over b}\log({1-q_{0}\over q_{0}}+2^{-b})=O({1\over b})$ Note that $1-\hat{p}^{1}(x_{m}|[u^{n}]_{b})={1\over n}\sum_{i=1}^{n}\mathds{1}_{[u_{i}]_{b}\neq x_{m}}$. Therefore, the optimization simplifies to a symbol-by-symbol optimization as follows:
${\hat{X}_{i}^{(1,b)}=\arg\min_{u\in\mathcal{U}}\Big{[}\log u^{2}+\frac{Y_{i}^{2}}{u^{2}}+\lambda(1+\gamma)\mathds{1}_{[u]_{b}\neq x_{m}}\Big{]}}.$
Note that if $[u]_{b}\neq x_{m}$, the loss function is minimized at ${\hat{x}}_{i}=|Y_{i}|$. Therefore, assuming that $b$ is large, to solve the optimization one needs to compare $\log Y_{i}^{2}+1+\lambda\gtrless\log x_{m}^{2}+{Y_{i}^{2}\over x_{m}^{2}}$, or
$1+\lambda\gtrless({Y_{i}\over x_{m}})^{2}-\log({Y_{i}\over x_{m}})^{2}.$ (4)
This optimization has the following closed-form solution
$\hat{X}_{i}^{(1,b)}=|Y_{i}|\mathds{1}_{|Y_{i}|/x_{m}\notin\mathcal{I}_{\lambda}}+x_{m}\mathds{1}_{|Y_{i}|/x_{m}\in\mathcal{I}_{\lambda}},$
where $\mathcal{I}_{\lambda}$ denotes an interval around one that is defined by (4).
3.2 Piecewise constant source
Consider a first order Markov process ${\bf X}$ such that given $X_{i}=x_{i}$, $X_{i+1}$ is distributed as $(1-q_{0})\delta_{x_{i}}+q_{0}{\rm Uniform}(x_{m},x_{M})$. Setting $k=2$ and again using the results of [27, Sec. 3.2], the BD-QMAP algorithm simplifies to
$\hat{X}^{n,(2,b)}=\arg\min_{u^{n}\in\mathcal{X}^{n}}\Big{[}{1\over n}\sum_{i=1}^{n}\big{(}\log u_{i}^{2}+\frac{Y_{i}^{2}}{u_{i}^{2}}\big{)}+(\lambda+{1\over b}\eta){N_{J}([u^{n}]_{b})\over n-1}\Big{]},$ (5)
where $\eta=-\lambda\log q_{0}-\log(1-q_{0}+q_{0}2^{-b})$ is a constant not depending on $u^{n}$. Here, $N_{J}(u^{n})=|\{i:u_{i}\neq u_{i+1}\}|$ denotes the number of jumps in $u^{n}$.
To provide a clearer understanding of how the BD-QMAP loss function operates, as well as the roles of its two key terms—the first ensuring reconstruction fidelity to the observations, and the second acting as a regularizer to promote the source structure—the following lemma offers an alternative representation of (5).
Lemma 1.
Solving the optimization in (5) is equivalent to solving the following optimization
$\hat{X}^{n,(2,b)}=\min_{k\in\{0,\ldots,n-1\}}\Big{(}\min_{(n_{1},\ldots,n_{k+1})}\sum_{j=1}^{k+1}n_{j}\log\big{(}{1\over n_{j}}\sum_{l\in\mathcal{I}_{j}}Y_{l}^{2}\big{)}+{n\over n-1}(\lambda+{1\over b})k\Big{)},$ (6)
where for any $k\in\{0,\ldots,n-1\}$, the inner optimization is over $(n_{1},\ldots,n_{k+1})\in(\mathds{N}^{+})^{k+1}$, such that $\sum_{i=1}^{k+1}n_{i}=n$. Moreover, given ${\bf n}=(n_{1},\ldots,n_{k+1})$, $\sum_{j=1}^{k+1}n_{j}=n$, for any $j\in\{1,\dots,k+1\}$,
$\mathcal{I}_{j}({\bf n})=\{\sum_{i=1}^{j-1}n_{i}+1,\ldots,\sum_{i=1}^{j}n_{i}\}.$ (7)
Moreover, for $i\in\mathcal{I}_{j}$,
$\hat{X}^{(2,b)}_{i}=\sqrt{{1\over n_{j}}\sum_{l\in\mathcal{I}_{j}}Y_{l}^{2}}.$ (8)
In other words, Lemma 1 implies that solving the BD-QMAP algorithm is equivalent to identifying the locations of the jumps such that, over each constant interval, the source input is estimated through appropriate averaging of the noisy observations, thereby minimizing the corresponding loss function. It is important to emphasize that if the regularization term enforcing the source structure were absent (i.e., if $\lambda=0$), the concavity of the $\log$ function would result in the solution equaling the absolute value of input noisy sequence, with the maximal number of jumps. On the other hand, choosing $\lambda$ too larges, will ensure that the output has no jumps.
The following theorem characterizes a lower bound on the performance achievable by BD-QMAP optimization (5).
Theorem 1.
Consider $X^{n}$ generated by a stationary first-order Markov source $\mathbf{X}=\{X_{i}\}_{i\geq 1}$, characterized by $p(X_{i+1}=x_{i+1}|X_{i}=x_{i})=(1-q_{0})\delta_{x_{i}}+q_{0}{\rm\pi_{c}}(x_{m},x_{M})$, where $\pi_{c}$ denotes the pdf of an absolutely continuous distribution with bounded support $[x_{m},x_{M}]$, $x_{m}>0$. Let $\operatorname{E}[X_{i}^{2}]=\eta^{2}$. Let ${\hat{X}}^{n}$ denote the solution of BD-QMAP optimization, when the number of jumps are known to be $k=k(n)$. Then,
$\operatorname{E}[{1\over n}\|X^{n}-{\hat{X}}^{n}\|_{2}^{2}]\geq$ $2q_{0}\eta^{2}(q_{0}(3-\sqrt{2/\pi}-\sqrt{2\pi})-q_{0}^{2}(2-\sqrt{2\pi})+\operatorname{E}[T(1-{1\over\sqrt{{\rm e}}}{(1-{1\over T})^{T\over 2}\over(1-{2\over T})^{T-1\over 2}})\mathbbmss{1}_{T\geq 3}])$
$-2M^{2}\Big{(}{\epsilon\over 1-\epsilon}+nq_{0}{\rm e}^{-\lfloor nq_{0}(1+\epsilon)\rfloor\left(q_{0}t_{1}+\ln\frac{1}{1+q_{0}t_{1}}\right)}+nq_{0}{\rm e}^{-\lfloor nq_{0}(1-\epsilon)\rfloor\left(q_{0}t_{2}+\ln\frac{1}{1+q_{0}t_{2}}\right)}\Big{)},$
where $T\sim\operatorname{Geometric}(q_{0})$ is representing the constant interval lengthes in the source.
Corollary 1.
Consider the same setup as in Theorem 1. Then,
$\operatorname{E}[{1\over n}\|X^{n}-{\hat{X}}^{n}\|_{2}^{2}]\geq$ $2q_{0}\eta^{2}(q_{0}(3-\sqrt{2/\pi}-\sqrt{2\pi})-q_{0}^{2}(2-\sqrt{2\pi})+\operatorname{E}[T(1-{1\over\sqrt{{\rm e}}}{(1-{1\over T})^{T\over 2}\over(1-{2\over T})^{T-1\over 2}})\mathbbmss{1}_{T\geq 3}])+\upsilon_{n},$
where $\upsilon_{n}=O(n^{-1/4})$.
4 Proofs
The following lemma will be used in the proof of the Theorem 1.
Lemma 2 (Concentration of Geometric [30]).
Let $T_{i}\overset{iid}{\sim}\operatorname{Geometric}(q_{0})\quad\forall i\in\{1,\dots,N\}$ , then for any $t>0$
$P\left(\left|\frac{\sum_{i=1}^{N}T_{i}}{N}-\frac{1}{q_{0}}\right|>t\right)\leq 2e^{-N\left(q_{0}t+\ln\frac{1}{1+q_{0}t}\right)}.$
Proof of Lemma 1.
Note that the set $\mathcal{U}^{n}$ can be partitioned into $n$ subsets depending on the number of jumps. That is, $\mathcal{U}^{n}=\bigcup_{i=0}^{n-1}\mathcal{U}_{k},$ where, for $k=0,\ldots,n-1$,
$\mathcal{U}_{k}=\{u^{n}\in\mathcal{U}^{n}:\;\sum_{i=2}^{n}\mathbbmss{1}_{u_{i}\neq u_{i-1}}=k\}.$
Then, the optimization in (5) can also be solved as follows
$\min_{k\in\{0,\ldots,n-1\}}\min_{u^{n}\in\mathcal{U}_{k}}[{1\over n}\sum_{j=1}^{n}(\log u_{i}^{2}+{Y_{i}^{2}\over u_{i}^{2}})+{1\over n-1}(\lambda+{1\over b}\eta)\sum_{i=2}^{n}\mathbbmss{1}_{u_{i}\neq u_{i-1}}]$
$\equiv\min_{k\in\{0,\ldots,n-1\}}\Big{(}\min_{u^{n}\in\mathcal{U}_{k}}\sum_{j=1}^{n}\big{(}\log u_{i}^{2}+{Y_{i}^{2}\over u_{i}^{2}}\big{)}+{n\over n-1}(\lambda+{1\over b})k\Big{)}.$ (9)
First, consider the inner optimization which is over all sequences with exactly $k$ jumps:
$\min_{u^{n}\in\mathcal{U}_{k}}\big{(}\log u_{i}^{2}+{Y_{i}^{2}\over u_{i}^{2}}\big{)}.$
Consider a sequence with constant intervals of length $n_{1},\ldots,n_{k+1}$ with corresponding values $a_{1},\ldots,a_{k+1}$. (Clearly $\sum_{i=1}^{k+1}n_{i}=n$.) That is, $u^{n}$ with $k$ jumps is written as
$u^{n}=\underbrace{a_{1},\ldots,a_{1}}_{n_{1}},\underbrace{a_{2},\ldots,a_{2}}_{n_{2}},\ldots,\underbrace{a_{k+1},\ldots,a_{k+1}}_{n_{k+1}}.$
Then, the optimization can be written as
$\min_{(n_{1},\ldots,n_{k+1}),(a_{1},\ldots,a_{k+1})\in\mathcal{U}^{k+1}}\big{(}\sum_{j=1}^{k+1}(n_{j}\log a_{i}^{2}+{1\over a_{j}^{2}}\sum_{l\in\mathcal{I}_{j}}Y_{l}^{2})\big{)}.$ (10)
where $\mathcal{I}_{j}$ is defined as (7). Then, it is easy to see that fixing the intervals, the optimal values can be found as
$\hat{a}_{j}^{2}={1\over n_{j}}\sum_{l\in\mathcal{I}_{j}}Y_{l}^{2}.$
Using this observation, the optimization in (10) can be written as
$\min_{(n_{1},\ldots,n_{k+1})}\sum_{j=1}^{k+1}n_{j}\log\big{(}{1\over n_{j}}\sum_{l\in\mathcal{I}_{j}}Y_{l}^{2}\big{)},$ (11)
which yields the desired result. ∎
Proof of Theorem 1.
Assuming that $X^{n}$ contains $k=k(n)$ jumps, it can be written as
$X^{n}=\underbrace{\alpha_{1},\ldots,\alpha_{1}}_{T_{1}},\underbrace{\alpha_{2},\ldots,\alpha_{2}}_{T_{2}},\ldots,\underbrace{\alpha_{k+1},\ldots,\alpha_{k+1}}_{T_{k+1}}.$
Moreover, when the number of jumps ($k$) is known apriori by the algorithm, (6) simplifies to
${\hat{X}}^{n}=\min_{(n_{1},\ldots,n_{k+1}):\;\sum_{i=1}^{k+1}n_{i}=n}\sum_{j=1}^{k+1}n_{j}\log\big{(}{1\over n_{j}}\sum_{l\in\mathcal{I}_{j}}Y_{l}^{2}\big{)}.$ (12)
Clearly, the expected error is minimized when the locations of the jumps are detected correctly (i.e., $n_{i}=T_{i}$, $i=1,\ldots,n+1$). We refer to this solution as the maximum likelihood (ML) solution, as it coincides with the ML solution when the number of jumps and their locations are known. As we just argued,
$\operatorname{E}[{1\over n}\|X^{n}-{\hat{X}}^{n}\|_{2}^{2}]\geq\operatorname{E}[{1\over n}\|X^{n}-{\hat{X}}^{n,{\rm(ML)}}\|_{2}^{2}].$
For $j=1,\ldots,k(n)+1$, define
$\hat{\alpha}_{j}=\sqrt{{1\over T_{j}}\sum_{l\in\mathcal{I}_{j}}Y_{l}^{2}}=\alpha_{j}\sqrt{{1\over T_{j}}\sum_{l\in\mathcal{I}_{j}}W_{l}^{2}}.$ (13)
Then, for $i\in\mathcal{I}_{j}$,
$\hat{X}_{i}^{\rm(ML)}=\hat{\alpha}_{j},$ (14)
which implies that
${1\over n}\|X^{n}-{\hat{X}}^{n,{\rm(ML)}}\|_{2}^{2}$ $={1\over n}\sum_{j=1}^{k}\alpha_{j}^{2}T_{j}\left(1-\sqrt{{1\over T_{j}}\sum_{l\in\mathcal{I}_{j}}W_{l}^{2}}\right)^{2}+T_{k+1}{\alpha_{k+1}^{2}\over n}\left(1-\sqrt{{1\over T_{k+1}}\sum_{l\in\mathcal{I}_{k+1}}W_{l}^{2}}\right)^{2}$
$\geq{1\over n}\sum_{j=1}^{k}\alpha_{j}^{2}T_{j}\left(1-\sqrt{{1\over T_{j}}\sum_{l\in\mathcal{I}_{j}}W_{l}^{2}}\right)^{2}$
$={q_{0}\over k}\sum_{j=1}^{k}\alpha_{j}^{2}T_{j}\left(1-\sqrt{{1\over T_{j}}\sum_{l\in\mathcal{I}_{j}}W_{l}^{2}}\right)^{2}+({k\over n}-q_{0}){1\over k}\sum_{j=1}^{k}\alpha_{j}^{2}T_{j}\left(1-\sqrt{{1\over T_{j}}\sum_{l\in\mathcal{I}_{j}}W_{l}^{2}}\right)^{2}$
$\stackrel{{\scriptstyle\rm(a)}}{{\geq}}{q_{0}\over k}\sum_{j=1}^{k}\alpha_{j}^{2}T_{j}\left(1-\sqrt{{1\over T_{j}}\sum_{l\in\mathcal{I}_{j}}W_{l}^{2}}\right)^{2}-|{k\over n}-q_{0}|{M^{2}\over k}\sum_{j=1}^{k}T_{j}(1+{1\over T_{j}}\sum_{l\in\mathcal{I}_{j}}W_{l}^{2})$
$\stackrel{{\scriptstyle\rm(b)}}{{\geq}}{q_{0}\over k}\sum_{j=1}^{k}\alpha_{j}^{2}T_{j}\left(1-\sqrt{{1\over T_{j}}\sum_{l\in\mathcal{I}_{j}}W_{l}^{2}}\right)^{2}-|{1\over q_{0}}-{n\over k}|M^{2}q_{0}(1+{1\over n}\sum_{l=1}^{n}W_{l}^{2}),$ (15)
where (a) follows because $(a-b)^{2}\leq a^{2}+b^{2}$ for all $a,b$ and (b) holds because $\sum_{j=1}^{k}T_{j}\leq n$ and $\sum_{j=1}^{k}\sum_{l\in\mathcal{I}_{j}}W_{l}^{2}\leq\sum_{l=1}^{n}W_{l}^{2}$.
Taking the expected value of both sides of (15), it follows that
$\operatorname{E}[{1\over n}\|X^{n}-{\hat{X}}^{n,{\rm(ML)}}\|_{2}^{2}]$ $\geq q_{0}\eta^{2}\operatorname{E}\left[T\Big{(}1-\sqrt{{1\over T}\sum_{i=1}^{T}W_{i}^{2}}\;\Big{)}^{2}\right]-M^{2}q_{0}\operatorname{E}\left[|{1\over q_{0}}-{n\over k}|(1+{1\over n}\sum_{l=1}^{n}W_{l}^{2})\right]$
$=q_{0}\eta^{2}\operatorname{E}\left[T\Big{(}1-\sqrt{{1\over T}\sum_{i=1}^{T}W_{i}^{2}}\;\Big{)}^{2}\right]-2M^{2}q_{0}\operatorname{E}\left[|{1\over q_{0}}-{n\over k}|\right],$ (16)
where the last line follows because the source $\mathbb{X}$ and the speckle noise process are independent. We first focus on the first term. Note that
$\operatorname{E}\left[T\Big{(}1-\sqrt{{1\over T}\sum_{j=1}^{T}w_{j}^{2}}\;\Big{)}^{2}\right]$ $=\operatorname{E}\left[T\Big{(}1+{1\over T}\sum_{j=1}^{T}w_{j}^{2}-2\sqrt{{1\over T_{i}}\sum_{j=1}^{T_{i}}w_{j}^{2}}\Big{)}\right]$
$=\operatorname{E}\left[2T-2T\sqrt{{1\over T}\sum_{j=1}^{T}w_{j}^{2}}\;\right]$
$=2\operatorname{E}\left[T-\sqrt{2T}{\Gamma({T+1\over 2})\over\Gamma({T\over 2})}\right],$ (17)
where the last line follows because
$\operatorname{E}\left[\sqrt{T\sum_{j=1}^{T}w_{j}^{2}}\;\right]$ $=\operatorname{E}\left[\operatorname{E}\left[\sqrt{T\sum_{j=1}^{T}w_{j}^{2}}\;\bigg{|}\;T\right]\right]=\operatorname{E}\left[\sqrt{2T}{\Gamma({T+1\over 2})\over\Gamma({T\over 2})}\right].$ (18)
It can be shown that, for all $x>0$,
$\sqrt{2\pi}{x}^{x+0.5}{\rm e}^{-x}{\rm e}^{{1\over 1+12x}}\leq\Gamma(x+1)\leq\sqrt{2\pi}{x}^{x+0.5}{\rm e}^{-x}{\rm e}^{{1\over 12x}}.$ (19)
Therefore, for $T>2$,
${\Gamma({T+1\over 2})\over\Gamma({T\over 2})}\;\leq\;$ ${{({T-1\over 2}})^{T\over 2}{\rm e}^{-(T-1)/2}{\rm e}^{{1\over 6(T-1)}}\over{({T\over 2}-1})^{T-1\over 2}{\rm e}^{-T/2+1}{\rm e}^{{1\over 6T-11}}}$
[…]
Therefore, we can bound (17) as
$2\operatorname{E}\left[T-\sqrt{2T}{\Gamma({T+1\over 2})\over\Gamma({T\over 2})}\right]$ $\geq 2q_{0}(1-{\sqrt{2}\over\Gamma(0.5)})+2q_{0}(1-q_{0})(2-2\Gamma(1.5))+2\operatorname{E}[T(1-{1\over\sqrt{{\rm e}}}{(1-{1\over T})^{T\over 2}\over(1-{2\over T})^{T-1\over 2}})\mathbbmss{1}_{T\geq 3}]$ (21)
Combining (16) and (21), it follows that
$\operatorname{E}[{1\over n}\|X^{n}-{\hat{X}}^{n,{\rm(ML)}}\|_{2}^{2}]\geq q_{0}$ $\eta^{2}(2q_{0}(1-{\sqrt{2}\over\Gamma(0.5)})+2q_{0}(1-q_{0})(2-2\Gamma(1.5))$ (22)
$+2\operatorname{E}[T(1-{1\over\sqrt{{\rm e}}}{(1-{1\over T})^{T\over 2}\over(1-{2\over T})^{T-1\over 2}})\mathbbmss{1}_{T\geq 3}])-2M^{2}q_{0}\operatorname{E}\left[|{1\over q_{0}}-{n\over k}|\right].$ (23)
To bound the last term, note that $k$ is a random variable that depends on $n$. Given $\epsilon>0$, define event $\mathcal{E}$ as
=$\mathcal{E}=\{nq_{0}(1-\epsilon)\leq k(n)\leq nq_{0}(1+\epsilon)\}.$ (24)
Note that conditioned on $\mathcal{E}$, we have $1-\epsilon\leq{k(n)\over nq_{0}}\leq 1+\epsilon$, which implies that
${1\over q_{0}(1+\epsilon)}-{1\over q_{0}}\leq{n\over k}-{1\over q_{0}}\leq{1\over q_{0}(1-\epsilon)}-{1\over q_{0}}.$
Using this observation, we can bound the remaining term in (23), as follows:
$\operatorname{E}\left[|{1\over q_{0}}-{n\over k}|\right]$ $=\operatorname{E}\left[|{1\over q_{0}}-{n\over k}|(\mathbbmss{1}_{\mathcal{E}}+\mathbbmss{1}_{\mathcal{E}^{c}})\right]$
$\leq{\epsilon\over q_{o}(1-\epsilon)}+nP(\mathcal{E}^{c}).$ (25)
Finally, to bound $P(\mathcal{E}^{c})$, note that
$P(k(n)\geq nq_{0}(1+\epsilon))$ $=P(\sum_{i=1}^{\lfloor nq_{0}(1+\epsilon)\rfloor}T_{i}<n)$
$=P\left({1\over\lfloor nq_{0}(1+\epsilon)\rfloor}\sum_{i=1}^{\lfloor nq_{0}(1+\epsilon)\rfloor}T_{i}-{1\over q_{0}}<{n\over\lfloor nq_{0}(1+\epsilon)\rfloor}-{1\over q_{0}}\right)$
$\leq P\left(|{1\over q_{0}}-{1\over\lfloor nq_{0}(1+\epsilon)\rfloor}\sum_{i=1}^{\lfloor nq_{0}(1+\epsilon)\rfloor}T_{i}|>{1\over q_{0}}-{n\over\lfloor nq_{0}(1+\epsilon)\rfloor}\right)$
$\leq{\rm e}^{-\lfloor nq_{0}(1+\epsilon)\rfloor\left(q_{0}t_{1}+\ln\frac{1}{1+q_{0}t_{1}}\right)},$ (26)
where the last line follows from Lemma 2, with
$t_{1}={1\over q_{0}}-{n\over\lfloor nq_{0}(1+\epsilon)\rfloor}.$
Similarly, again using Lemma 2, we have
$P(k(n)\leq nq_{0}(1-\epsilon))$ $=P(\sum_{i=1}^{\lfloor nq_{0}(1-\epsilon)\rfloor}T_{i}>n)$
$=P({1\over\lfloor nq_{0}(1-\epsilon)\rfloor}\sum_{i=1}^{\lfloor nq_{0}(1-\epsilon)\rfloor}T_{i}-{1\over q_{0}}>{n\over\lfloor nq_{0}(1-\epsilon)\rfloor}-{1\over q_{0}})$
$\leq{\rm e}^{-\lfloor nq_{0}(1-\epsilon)\rfloor\left(q_{0}t_{2}+\ln\frac{1}{1+q_{0}t_{2}}\right)},$ (27)
with
$t_{2}={n\over\lfloor nq_{0}(1-\epsilon)\rfloor}-{1\over q_{0}}.$
Combining (25), (26) and (27) with (23) yields the desired result. That is,
$\operatorname{E}[{1\over n}\|X^{n}-{\hat{X}}^{n}\|_{2}^{2}]\geq$ $q_{0}\eta^{2}(2q_{0}(1-{\sqrt{2}\over\Gamma(0.5)})+2q_{0}(1-q_{0})(2-2\Gamma(1.5))+2\operatorname{E}[T(1-{1\over\sqrt{{\rm e}}}{(1-{1\over T})^{T\over 2}\over(1-{2\over T})^{T-1\over 2}})\mathbbmss{1}_{T\geq 3}])$ (28)
$-2M^{2}q_{0}\Big{(}{\epsilon\over q_{o}(1-\epsilon)}+n{\rm e}^{-\lfloor nq_{0}(1+\epsilon)\rfloor\left(q_{0}t_{1}+\ln\frac{1}{1+q_{0}t_{1}}\right)}+n{\rm e}^{-\lfloor nq_{0}(1-\epsilon)\rfloor\left(q_{0}t_{2}+\ln\frac{1}{1+q_{0}t_{2}}\right)}\Big{)}.$ (29)
Proof of Corollary 1.
For $x\geq 0$ we can show [31] that $\ln(1+x)\leq x\frac{x+6}{2x+6}$ thus,
$x-\ln(1+x)\geq\frac{x^{2}}{2x+6}.$ (30)
In addition, we have
$1-\frac{x}{\lfloor x(1+x^{-1/4})\rfloor}=\Omega(\frac{1}{x^{1/3}}),$ (31)
and,
$\frac{x}{\lfloor x(1-x^{-1/4})\rfloor}-1=\Omega(\frac{1}{x^{1/3}}).$ (32)
(31) holds for $x\geq x_{0}$,
$1-\frac{x}{\lfloor x+x^{3/4}\rfloor}$ $=1-\frac{x}{\lfloor x+x^{3/4}\rfloor-x-x^{3/4}+x+x^{3/4}}$
$\geq 1-\frac{x}{-1+x+x^{3/4}}$
$=\frac{x^{3/4}-1}{x+x^{3/4}-1}$
$=\frac{1}{x^{1/3}}\underbrace{\left(\frac{x^{3/4}-1}{x^{2/3}+x^{5/12}-x^{-1/3}}\right)}_{(*)}$
$\geq\frac{1}{x^{1/3}},$
where $x_{0}$ is some point after which ()(*)( ∗ ) remains greater than one. Similar steps proves (32).
Let $\epsilon=(nq_{0})^{-1/4}$ in the result of Theorem 1, using (31), (32) we have $q_{0}t_{1}=\Omega(n^{-1/3})$ and $q_{0}t_{2}=\Omega(n^{-1/3})$, and applying (30) we have
$\operatorname{E}[{1\over n}\|X^{n}-{\hat{X}}^{n}\|_{2}^{2}]\geq$ $q_{0}\eta^{2}(2q_{0}(1-{\sqrt{2}\over\Gamma(0.5)})+2q_{0}(1-q_{0})(2-2\Gamma(1.5))+2\operatorname{E}[T(1-{1\over\sqrt{{\rm e}}}{(1-{1\over T})^{T\over 2}\over(1-{2\over T})^{T-1\over 2}})\mathbbmss{1}_{T\geq 3}])$
$-2M^{2}\Big{(}{(nq_{0})^{-1/4}\over 1-(nq_{0})^{-1/4}}+q_{0}n{\rm e}^{-\lfloor nq_{0}(1+(nq_{0})^{-1\over 4})\rfloor\left(c_{1}^{2}n^{-2\over 3}\over 2n^{-1\over 3}+6\right)}+q_{0}n{\rm e}^{-\lfloor nq_{0}(1-(nq_{0})^{-1\over 4})\rfloor\left(c_{2}^{2}n^{-2\over 3}\over 2n^{-1\over 3}+6\right)}\Big{)},$
where $c_{1},c_{2}\in\mathds{R}^{+}$ are constants. Second line converges to zero as $n\rightarrow\infty$ with $O(n^{-1/4})$. ∎
5 Numerical experiments
Figure 1: MSE lower bound of the genie-aided ML despeckler as derived in Theorem 1 and the proposed BD-QMAP despecker for piecewise constant source with parameter $q_{0}$. ($\operatorname{PSNR}=10\log_{10}(\operatorname{MAX}^{2}/\operatorname{MSE})$ where $\operatorname{MAX}$ is the maximum value of signal’s support set)
In the following, we first address the practical implementation of the BD-QMAP optimization problem. We then compare the performance of BD-QMAP with the lower bound derived in Theorem 1. Finally, we perform comparative studies on despeckling for piecewise-constant sources and extending the BD-QMAP to be applied on high dimensional images.
5.1 Algorithmic considerations
As noted in Remark 1, addressing the distinct search spaces of the log-likelihood and regularizer in BD-QMAP is essential for practical implementation, and limiting the search to the quantized space ensures feasibility. This restriction enables the discrete suboptimal optimization to be efficiently performed using the Viterbi algorithm [32]. Therefore, to tackle the optimization problem of BD-QMAP, we begin by limiting the search space in (3) to $b$-bit quantized realizations, denoted as $\mathcal{X}^{n}_{b}$. We call this suboptimal despeckler BD-QMAPb, as it is optimizing the both terms of (3) in the $b$-bit quantized space. In other words, BD-QMAPb solves the following optimization
$\operatorname*{arg\,min}_{u^{n}\in\mathcal{X}^{n}_{b}}{1\over n}\sum_{i=k}^{n}(\log u_{i}^{2}+\frac{y_{i}^{2}}{u_{i}^{2}}+\frac{\lambda}{b}w_{u_{i-k+1}^{i}}),$ (33)
where $c_{\mathbf{w}}(u^{n})$ is scaled by $(n-k+1)/n$ without an impact on the optimal solution as it can be absorbed in the hyper-parameter $\lambda$. Since similar to the the additive case, the negative log likelihood function of $u^{n}$ is the sum of $n$ element-wise distinct functions, with the $i$-th function only depending on $u_{i}$, the Viterbi algorithm [32] can still be employed to solve (33).
The Viterbi algorithm operates on a trellis diagram with $n$ stages, each having $|\mathcal{X}_{b}|^{k}$ states. State $s_{i}$ in the diagram corresponds to the subsequence $u_{i-k+1}^{i}$ of length $k$ in the candidate solution. Let the cost associated to the state $s_{i}=u^{i}_{i-k+1}$ in the diagram follows the recursive rule as,
$C_{i}(s_{i})=\min_{s^{\prime}\in\mathcal{S}(s_{i})}(\log u_{i}^{2}+\frac{y_{i}^{2}}{u_{i}^{2}}+\frac{\lambda}{b}w_{u^{i}_{i-k+1}})+C_{i-1}(s^{\prime}),$
for any $i\in\{k,\dots,n\}$, where $\mathcal{S}(s_{i})$ denotes the set of all closest states that transition into the state $s_{i}$ and $|\mathcal{S}(s_{i})|=|\mathcal{X}_{b}|$. The state $s^{\prime}=u_{i-k}^{i-1}$ represents the previous state leading to $s_{i}$. Thus BD-QMAPb solution can be found recursively starting minimizing $C_{n}(s_{n})$.
5.2 Theoretical lower bound
In Theorem 1, a lower bound on the best achievable reconstruction error was derived for the piecewise constant source. To explore this bound in practice, we compare it to an ideal despeckler for piecewise constant sources, referred to as the genie-aided maximum likelihood (ML) despeckler. This estimator has access to the exact jump locations in the ground truth data in addition to the noisy observations. With such side information, the reconstruction error is minimized using Equation 8 for each constant segment between consecutive jump locations. For comparison, we evaluated the performance of this genie-aided despeckler against BD-QMAP. The BD-QMAP implementation, as detailed in Section 5.3.1, was applied to piecewise constant sources with $q_{0}\in\{0.001,0.01,0.05,0.1\}$. Each data point in the figure represents the average error over a test set of 100 realizations with fixed $q_{0}$, each of length $n=10^{6}$.
5.3 Comparative analysis of despeckling for piecewise constant sources
Figure 2: Piecewise constant source ($X$) with parameter $q_{0}=0.001$ sampled under speckle noise ($Y$), enhanced Lee and BD-QMAPb despeckled reconstructions, as discussed in Section 5.
For numerical experiments with the piecewise constant source introduced in Section 3.2 we assume the absolute continous distribution to be uniform on the interval $[0,1]$. A sample realization of this source under speckle noise is shown in Figure 2. Considering this source, the pivotal parameter which directly affects the performance of the despeckler is $q_{0}$, i.e., the probability of signal level-change which also determines the level of structure for the analog source [29]. To learn the weights of the Equation 2 to be used in the regularizer term of the BD-QMAP loss function, for a fixed $q_{0}$, we sample the random process for $10^{7}$ points and find the empirical second order distribution by counting the occurances of the $b$ bit quantized samples of length two. Key advantages of BD-QMAP over other despecklers in the literature in practice are two fold. First, the BD-QMAP assumes no upperbound on the speckle power level on the observation as opposed to [13, 14], and also [8] in which the filter is bypassed completely in high noise level regimes. Second, the BD-QMAP method is applicable to any given source without requiring restrictive assumptions on the signal’s prior distribution. This flexibility contrasts with the strong assumptions imposed by other Bayesian filters, such as those in [7, 10, 6]. Each entry in the Table 1 is the despeckling performance for a test dataset of size $100$ of piecewise constant signals of length $n=10^{6}$.
5.3.1 BD-QMAP implementation
Using the Viterbi implementation described in Section 5.1 for $k=2$, we can efficiently implement BD-QMAPb. As noted in Remark 1, the fidelity term in BD-QMAP optimization is not restricted to a quantized space. Therefore, to enhance the BD-QMAPb solution, we refine it by identifying the detected jump locations and computing the ML estimates on the intervals between them using Equation 8. This refinement significantly improves performance, especially for highly structured sources (i.e., when $q_{0}$ is small). As shown in Table 1, the despeckling performance of BD-QMAPb for highly structured sources improves substantially when the initial estimate is refined for the final BD-QMAP solution. Notably, even a low-complexity implementation with $b=2$ is effective in detecting jump locations, with the refinement step yielding significant gains.
5.3.2 Regularization and quantization in BD-QMAP
To explore the effectiveness of the regularizer of the BD-QMAP in finding the structure of the source, and yeilding lower errors, we conducted experiments for a range of values for $\lambda$ in Equation 3, as shown in Figure 3. It can be seen that when $\lambda=0$ the despeckler is not able to distinguish between different levels of structuredness (as measured by $q_{0}$), however a small regularization would let the denoiser to recover the empirical distribution of the ground-truth distribution in the reconstruction. Finally, as expected, a large value of $\lambda$ would decrease the fidelity of the reconstruction to the observed noisy signal. The number of bits used to quantize the source for the learned weights significantly impacts both the final performance and the time complexity of the optimization problem in Equation 3, as the size of the search space increases exponentially with the number of bits.
In general, achieving better reconstruction results necessitates selecting the number of bits $b$ in BD-QMAP as high as computationally feasible. However, for highly structured data (e.g., high-resolution images), a relatively small number of bits can suffice to uncover the structure of the source from noisy observations. As demonstrated in Figure 3, when $q_{0}=0.001$ (indicating a highly structured source), the suboptimal solution of BD-QMAPb with just 2 bits effectively captures the source structure. Upon refining the solution to account for quantization error, it achieves a higher PSNR than BD-QMAPb with 3 bits. This experiment suggests that for highly structured sources, selecting fewer bits is often sufficient to match the distribution, avoiding the computational challenges associated with higher bit counts.
Table 1: Despeckling Performance Comparison (PSNR) for Piecewise Constant Source.
$q_{0}=0.1$ $q_{0}=0.01$ $q_{0}=0.001$
Speckled Source 8.710 8.742 9.055
Box Car Filter 13.226 16.321 16.976
Frost Filter [33] 12.923 13.900 14.226
Total Variation [14] 9.339 10.895 11.483
Lee Filter [6] / Enhanced [8] 10.505 / 13.865 18.402 / 19.703 22.303 / 22.577
Kuan Filter [7] / Enhanced [8] 11.618 / 14.041 19.650 / 20.375 23.017 / 23.244
BD-QMAPb (b=2 / b=3) 14.786 / 15.478 17.909 / 21.613 19.530 / 24.363
BD-QMAP (b=2 / b=3) 14.696 / 15.072 20.961 / 22.518 27.052 / 30.831
Figure 3: Effect of hyperparameter $\lambda$ in BD-QMAPb and BD-QMAP optimization for different choices of bits ($b$) and source structure ($q_{0}$).
5.3.3 Linear adaptive filters
All the linear adaptive filters in Table 1 require a window size hyper-parameter that defines the area over which the local statistics will be calculated. In these experiments this value is set to $1/2q_{0}$, half of the expected interval lengths of constant pieces in the source.
We also conduct experiments with an Enhanced version of the filters which is based on the idea of thresholding local variance, as suggested in [8]. To improve the final performance of the filters, three statiscical classes are considered which are known as heterogeneity adjustments in the literature. The simple modification is to distinguish between three region of operation: homogeneous, heterogenous, and strongly speckled. To decide in which region the filter operates, if the observation variation coefficient defined as $C_{Y}\triangleq\frac{\operatorname{Var}(Y)}{\operatorname{E}[Y]}$ of the local window falls below a threshold of noise variation coefficient, similarly defined to be $C_{W}\triangleq\frac{\operatorname{Var}(W)}{\operatorname{E}[W]}$, we only do keep the mean window mean value as the despeckled pixel. On the other hand, if $C_{Y}\geq C_{\max}$, the filter output is bypassed and the despeckled pixel is equal to the speckled observed pixel, where empirically the optimal value of $C_{\max}$ is set to $\sqrt{3}C_{W}$ [8]. The Enhanced version of the filters mentioned in Table 1 refer to applying the filter only when $C_{W}\leq C_{Y}\leq C_{\max}$.
6 Conclusion
A novel Bayesian despeckling method is proposed for structured sources. Its performance is analyzed for piecewise constant structured sources, and a lower bound on the minimum achievable MSE is derived. The results demonstrate that the despeckler effectively detects edges under speckle noise, achieving superior performance compared to other theoretically grounded approaches.
Acknowledgment
AZ and SJ were supported by NSF grant CCF-2237538.

### v2

@@
 These include the use of generative adversarial networks as demonstrated in [25] and diffusion-based probabilistic models as explored in [26].
-1.3 Our contribution
+1.3 Our contributions
 We propose Bayesian Despeckling via QMAP (BD-QMAP), a novel despeckling algorithm inspired by the Quantized Maximum a Posteriori (Q-MAP) estimator [27], originally developed for additive noise.
@@
 Then, given weights $\mathbf{w}=(w_{a^{k}}:a^{k}\in\mathcal{U}_{b}^{k})$, the weight assigned to sequence $u^{n}\in\mathcal{U}^{n}$ is defined as
-$c_{\mathbf{w}}(u^{n})=\sum_{a^{k}\in\mathcal{X}^{k}_{b}}w_{a^{k}}\hat{p}^{k}(a^{k}|u^{n})(a^{k}|[X^{k}]_{b}).$ (2)
+$c_{\mathbf{w}}(u^{n})=\sum_{a^{k}\in\mathcal{X}^{k}_{b}}w_{a^{k}}\hat{p}^{k}(a^{k}|u^{n}).$ (2)
 Finally, the BD-QMAP method, recovers $X^{n}$ as $\hat{X}^{n,(k,b)}$ defined as
@@
 The first term represents a fidelity criterion derived from the negative likelihood of the observations given the signal under the multiplicative noise model.
-Note that given $X^{n}=x^{n}$, $Y^{n}\sim\mathcal{N}({\bf 0},X^{2})$, where $X=\mathop{\rm Diag}\nolimits({\bf x})$.
+Note that given $X^{n}=x^{n}$, $Y^{n}\sim\mathcal{N}({\bf 0},X^{2})$, where $X=\mathop{\rm Diag}\nolimits(x^{n})$.
 Therefore,
@@
 In other words, the weights are expected to summarize the source’s $n$-dimensional distribution characterized by $p(x^{n})$ into a finite number $|\mathcal{X}_{b}|^{k}$ of positive weights.
-A key advantage of this approach is that it leads to a tractable approach for modeling and utilizing the sources structure.
+A key advantage of this approach is that it leads to a tractable approach for modeling and utilizing the source’s structure.
 It can be shown that for structured sources, i.e., the sources with information dimension strictly less than one [28, 29], identified by having singularities in their distributions, e.g., spike and slap distribution, the weights ${\bf w}$ can be divided into two groups: 1) A small set of weights that identify the key structure of the source, and 2) a large set of weights that have negligible impact in the optimization [27].
@@
 To address this complexity, the optimization can be simplified by restricting the search space to $\mathcal{X}_{b}^{n}$, the space of quantized sequences.
-As explained in Section 5.1, this restriction enables the application of the Viterbi algorithm for efficient optimization.
+As will be explained in Section 5.1, this restriction enables the application of the Viterbi algorithm for efficient optimization.
 While this approach is suboptimal compared to solving the original problem, it offers significant computational advantages.
@@
 3.1 Structured memoryless source
-Consider an i.i.d process ${\bf X}$ such that $X_{i}\sim(1-q_{0})\delta_{x_{m}}+q_{0}{\rm Uniform}(x_{m},x_{M})$, Setting $k=1$ and using the results of [27, Sec.
-3.1] to simplfy the second term in (3), the BD-QMAP algorithm can be written as
+Consider an i.i.d.
+process ${\bf X}$ such that $X_{i}\sim(1-q_{0})\delta_{x_{m}}+q_{0}{\rm Uniform}(x_{m},x_{M})$, Setting $k=1$ and using the results of [27, Sec.
+3.1] to simplify the second term in (3), the BD-QMAP algorithm can be written as
 ${\hat{X}^{n,(1,b)}=\arg\min_{u^{n}\in\mathcal{U}^{n}}\Big{[}{1\over n}\sum_{i=1}^{n}\big{(}\log u_{i}^{2}+\frac{Y_{i}^{2}}{u_{i}^{2}}\big{)}+\lambda(1+\gamma)(1-\hat{p}^{1}(x_{m}|[u^{n}]_{b}))\Big{]}},$
@@
 To provide a clearer understanding of how the BD-QMAP loss function operates, as well as the roles of its two key terms—the first ensuring reconstruction fidelity to the observations, and the second acting as a regularizer to promote the source structure—the following lemma offers an alternative representation of (5).
+Consider ${\bf n}=(n_{1},\ldots,n_{k+1})\in(\mathbb{N}^{+})^{k+1}$, such that $\sum_{j=1}^{k+1}n_{j}=n$. Then, for any $j\in\{1,\dots,k+1\}$, define
+$\mathcal{I}_{j}({\bf n})=\Big{\{}\sum_{i=1}^{j-1}n_{i}+1,\ldots,\sum_{i=1}^{j}n_{i}\Big{\}}.$ (6)
 Lemma 1.
@@
 $\hat{X}^{n,(2,b)}=\min_{k\in\{0,\ldots,n-1\}}\Big{(}\min_{(n_{1},\ldots,n_{k+1})}\sum_{j=1}^{k+1}n_{j}\log\big{(}{1\over n_{j}}\sum_{l\in\mathcal{I}_{j}}Y_{l}^{2}\big{)}+{n\over n-1}(\lambda+{1\over b})k\Big{)},$ (7)
-where for any $k\in\{0,\ldots,n-1\}$, the inner optimization is over $(n_{1},\ldots,n_{k+1})\in(\mathds{N}^{+})^{k+1}$, such that $\sum_{i=1}^{k+1}n_{i}=n$.
-Moreover, given ${\bf n}=(n_{1},\ldots,n_{k+1})$, $\sum_{j=1}^{k+1}n_{j}=n$, for any $j\in\{1,\dots,k+1\}$,
-$\mathcal{I}_{j}({\bf n})=\{\sum_{i=1}^{j-1}n_{i}+1,\ldots,\sum_{i=1}^{j}n_{i}\}.$ (7)
-Moreover, for $i\in\mathcal{I}_{j}$,
+where for any $k\in\{0,\ldots,n-1\}$, the inner optimization is over $(n_{1},\ldots,n_{k+1})\in(\mathbb{N}^{+})^{k+1}$, such that $\sum_{i=1}^{k+1}n_{i}=n$.
+Moreover, for $i\in\mathcal{I}_{j}({\bf n})$,
 $\hat{X}^{(2,b)}_{i}=\sqrt{{1\over n_{j}}\sum_{l\in\mathcal{I}_{j}}Y_{l}^{2}}.$ (8)
@@
 It is important to emphasize that if the regularization term enforcing the source structure were absent (i.e., if $\lambda=0$), the concavity of the $\log$ function would result in the solution equaling the absolute value of input noisy sequence, with the maximal number of jumps.
-On the other hand, choosing $\lambda$ too larges, will ensure that the output has no jumps.
+On the other hand, choosing $\lambda$ too large, will ensure that the output has no jumps.
 The following theorem characterizes a lower bound on the performance achievable by BD-QMAP optimization (5).
@@
 Consider $X^{n}$ generated by a stationary first-order Markov source $\mathbf{X}=\{X_{i}\}_{i\geq 1}$, characterized by $p(X_{i+1}=x_{i+1}|X_{i}=x_{i})=(1-q_{0})\delta_{x_{i}}+q_{0}{\rm\pi_{c}}(x_{m},x_{M})$, where $\pi_{c}$ denotes the pdf of an absolutely continuous distribution with bounded support $[x_{m},x_{M}]$, $x_{m}>0$.
-Let $\operatorname{E}[X_{i}^{2}]=\eta^{2}$.
 Let ${\hat{X}}^{n}$ denote the solution of BD-QMAP optimization, when the number of jumps are known to be $k=k(n)$.
 Then,
-$\operatorname{E}[{1\over n}\|X^{n}-{\hat{X}}^{n}\|_{2}^{2}]\geq$ $2q_{0}\eta^{2}(q_{0}(3-\sqrt{2/\pi}-\sqrt{2\pi})-q_{0}^{2}(2-\sqrt{2\pi})+\operatorname{E}[T(1-{1\over\sqrt{{\rm e}}}{(1-{1\over T})^{T\over 2}\over(1-{2\over T})^{T-1\over 2}})\mathbbmss{1}_{T\geq 3}])$
-$-2M^{2}\Big{(}{\epsilon\over 1-\epsilon}+nq_{0}{\rm e}^{-\lfloor nq_{0}(1+\epsilon)\rfloor\left(q_{0}t_{1}+\ln\frac{1}{1+q_{0}t_{1}}\right)}+nq_{0}{\rm e}^{-\lfloor nq_{0}(1-\epsilon)\rfloor\left(q_{0}t_{2}+\ln\frac{1}{1+q_{0}t_{2}}\right)}\Big{)},$
-where $T\sim\operatorname{Geometric}(q_{0})$ is representing the constant interval lengthes in the source.
+$\operatorname{E}[{1\over n}\|X^{n}-{\hat{X}}^{n}\|_{2}^{2}]\geq 2q_{0}\eta^{2}\Big{(}q_{0}c_{1}+q_{0}^{2}c_{2}+\operatorname{E}[T(1-{1\over\sqrt{{\rm e}}}{(1-{1\over T})^{T\over 2}\over(1-{2\over T})^{T-1\over 2}})\mathds{1}_{T\geq 3}]\Big{)}-\upsilon_{n},$ (9)
+where $\eta^{2}=\operatorname{E}[X_{i}^{2}]$, $c_{1}=3-\sqrt{2/\pi}-2\sqrt{\pi}$ and $c_{2}=\sqrt{\pi}-2$, and $T\sim\operatorname{Geometric}(q_{0})$. Here,
+$\upsilon_{n}=2x_{M}^{2}\Big{(}{\epsilon\over 1-\epsilon}+nq_{0}{\rm e}^{-\lfloor nq_{0}(1+\epsilon)\rfloor\left(q_{0}t_{1}+\ln\frac{1}{1+q_{0}t_{1}}\right)}+nq_{0}{\rm e}^{-\lfloor nq_{0}(1-\epsilon)\rfloor\left(q_{0}t_{2}+\ln\frac{1}{1+q_{0}t_{2}}\right)}\Big{)},$
+with $t_{1}={1/q_{0}}-{n/\lfloor nq_{0}(1+\epsilon)\rfloor}$ and $t_{2}={n/\lfloor nq_{0}(1-\epsilon)\rfloor}-{1/q_{0}}$.
+The following corollary, presents a simplified form of Theorem 1, for large values of $n$.
 Corollary 1.
 Consider the same setup as in Theorem 1. Then,
-$\operatorname{E}[{1\over n}\|X^{n}-{\hat{X}}^{n}\|_{2}^{2}]\geq$ $2q_{0}\eta^{2}(q_{0}(3-\sqrt{2/\pi}-\sqrt{2\pi})-q_{0}^{2}(2-\sqrt{2\pi})+\operatorname{E}[T(1-{1\over\sqrt{{\rm e}}}{(1-{1\over T})^{T\over 2}\over(1-{2\over T})^{T-1\over 2}})\mathbbmss{1}_{T\geq 3}])+\upsilon_{n},$
+$\operatorname{E}[{1\over n}\|X^{n}-{\hat{X}}^{n}\|_{2}^{2}]\geq$ $2q_{0}\eta^{2}(q_{0}c_{1}-q_{0}^{2}c_{2}+\operatorname{E}[T(1-{1\over\sqrt{{\rm e}}}{(1-{1\over T})^{T\over 2}\over(1-{2\over T})^{T-1\over 2}})\mathds{1}_{T\geq 3}])-\upsilon_{n},$
 where $\upsilon_{n}=O(n^{-1/4})$.
+Finally, to better understand the term in (9) that involves expectation over $T$, note that for small values of $q_{0}$, with high probability, $T$ is large.
+On the other hand, for large values of $T$, $f(T)=1/4+O(T^{-1})$, as
+$f(T)=T(1-{1\over\sqrt{{\rm e}}}{(1-{1\over T})^{T\over 2}\over(1-{2\over T})^{T-1\over 2}})$ $=T(1-\exp(-\frac{1}{2}+\frac{T}{2}\ln(1-\frac{1}{T})-\frac{T-1}{2}\ln(1-\frac{2}{T})))$
+$=T(1-\exp(-\frac{1}{2}+\frac{T}{2}(-\frac{1}{T}-\frac{1}{2T^{2}}+O(T^{-3}))-\frac{T-1}{2}(-\frac{2}{T}-\frac{2}{T^{2}}+O(T^{-3}))))$
+$=T(1-\exp(-\frac{1}{4T}-\frac{1}{T^{2}}+O(T^{-2})))$
+$=T(1-1+\frac{1}{4T}+O(T^{-2}))$
+$=\frac{1}{4}+O(T^{-1}).$
 4 Proofs
@@
 Lemma 2 (Concentration of Geometric [30]).
-Let $T_{i}\overset{iid}{\sim}\operatorname{Geometric}(q_{0})\quad\forall i\in\{1,\dots,N\}$ , then for any $t>0$
-$P\left(\left|\frac{\sum_{i=1}^{N}T_{i}}{N}-\frac{1}{q_{0}}\right|>t\right)\leq 2e^{-N\left(q_{0}t+\ln\frac{1}{1+q_{0}t}\right)}.$
-Proof of Lemma 1.
-Note that the set $\mathcal{U}^{n}$ can be partitioned into $n$ subsets depending on the number of jumps. That is, $\mathcal{U}^{n}=\bigcup_{i=0}^{n-1}\mathcal{U}_{k},$ where, for $k=0,\ldots,n-1$,
-$\mathcal{U}_{k}=\{u^{n}\in\mathcal{U}^{n}:\;\sum_{i=2}^{n}\mathbbmss{1}_{u_{i}\neq u_{i-1}}=k\}.$
+Assume that $T_{i}\overset{iid}{\sim}\operatorname{Geometric}(q_{0})$, $i=1,\ldots,N$. Then for any $t>0$
+$\operatorname{P}\left(\left|\frac{\sum_{i=1}^{N}T_{i}}{N}-\frac{1}{q_{0}}\right|>t\right)\leq 2e^{-N\left(q_{0}t+\ln\frac{1}{1+q_{0}t}\right)}.$
+4.1 Proof of Lemma 1
+Note that the set $\mathcal{U}^{n}$ can be partitioned into $n$ subsets depending on the number of jumps. That is, $\mathcal{U}^{n}=\bigcup_{k=0}^{n-1}\mathcal{U}_{k},$ where, for $k=0,\ldots,n-1$,
+$\mathcal{U}_{k}=\{u^{n}\in\mathcal{U}^{n}:\;\sum_{i=2}^{n}\mathds{1}_{u_{i}\neq u_{i-1}}=k\}.$
 Then, the optimization in (5) can also be solved as follows
-$\min_{k\in\{0,\ldots,n-1\}}\min_{u^{n}\in\mathcal{U}_{k}}[{1\over n}\sum_{j=1}^{n}(\log u_{i}^{2}+{Y_{i}^{2}\over u_{i}^{2}})+{1\over n-1}(\lambda+{1\over b}\eta)\sum_{i=2}^{n}\mathbbmss{1}_{u_{i}\neq u_{i-1}}]$
+$\min_{k\in\{0,\ldots,n-1\}}\min_{u^{n}\in\mathcal{U}_{k}}[{1\over n}\sum_{j=1}^{n}(\log u_{i}^{2}+{Y_{i}^{2}\over u_{i}^{2}})+{1\over n-1}(\lambda+{1\over b}\eta)\sum_{i=2}^{n}\mathds{1}_{u_{i}\neq u_{i-1}}]$
 $\equiv\min_{k\in\{0,\ldots,n-1\}}\Big{(}\min_{u^{n}\in\mathcal{U}_{k}}\sum_{j=1}^{n}\big{(}\log u_{i}^{2}+{Y_{i}^{2}\over u_{i}^{2}}\big{)}+{n\over n-1}(\lambda+{1\over b})k\Big{)}.$ (10)
@@
 $\min_{(n_{1},\ldots,n_{k+1})}\sum_{j=1}^{k+1}n_{j}\log\big{(}{1\over n_{j}}\sum_{l\in\mathcal{I}_{j}}Y_{l}^{2}\big{)},$ (12)
-which yields the desired result. ∎
-Proof of Theorem 1.
+which yields the desired result.
+4.2 Proof of Theorem 1
 Assuming that $X^{n}$ contains $k=k(n)$ jumps, it can be written as
@@
 $={q_{0}\over k}\sum_{j=1}^{k}\alpha_{j}^{2}T_{j}\left(1-\sqrt{{1\over T_{j}}\sum_{l\in\mathcal{I}_{j}}W_{l}^{2}}\right)^{2}+({k\over n}-q_{0}){1\over k}\sum_{j=1}^{k}\alpha_{j}^{2}T_{j}\left(1-\sqrt{{1\over T_{j}}\sum_{l\in\mathcal{I}_{j}}W_{l}^{2}}\right)^{2}$
-$\stackrel{{\scriptstyle\rm(a)}}{{\geq}}{q_{0}\over k}\sum_{j=1}^{k}\alpha_{j}^{2}T_{j}\left(1-\sqrt{{1\over T_{j}}\sum_{l\in\mathcal{I}_{j}}W_{l}^{2}}\right)^{2}-|{k\over n}-q_{0}|{M^{2}\over k}\sum_{j=1}^{k}T_{j}(1+{1\over T_{j}}\sum_{l\in\mathcal{I}_{j}}W_{l}^{2})$
-$\stackrel{{\scriptstyle\rm(b)}}{{\geq}}{q_{0}\over k}\sum_{j=1}^{k}\alpha_{j}^{2}T_{j}\left(1-\sqrt{{1\over T_{j}}\sum_{l\in\mathcal{I}_{j}}W_{l}^{2}}\right)^{2}-|{1\over q_{0}}-{n\over k}|M^{2}q_{0}(1+{1\over n}\sum_{l=1}^{n}W_{l}^{2}),$ (15)
+$\stackrel{{\scriptstyle\rm(a)}}{{\geq}}{q_{0}\over k}\sum_{j=1}^{k}\alpha_{j}^{2}T_{j}\left(1-\sqrt{{1\over T_{j}}\sum_{l\in\mathcal{I}_{j}}W_{l}^{2}}\right)^{2}-|{k\over n}-q_{0}|{x_{M}^{2}\over k}\sum_{j=1}^{k}T_{j}(1+{1\over T_{j}}\sum_{l\in\mathcal{I}_{j}}W_{l}^{2})$
+$\stackrel{{\scriptstyle\rm(b)}}{{\geq}}{q_{0}\over k}\sum_{j=1}^{k}\alpha_{j}^{2}T_{j}\left(1-\sqrt{{1\over T_{j}}\sum_{l\in\mathcal{I}_{j}}W_{l}^{2}}\right)^{2}-|{1\over q_{0}}-{n\over k}|x_{M}^{2}q_{0}(1+{1\over n}\sum_{l=1}^{n}W_{l}^{2}),$ (16)
 where (a) follows because $(a-b)^{2}\leq a^{2}+b^{2}$ for all $a,b$ and (b) holds because $\sum_{j=1}^{k}T_{j}\leq n$ and $\sum_{j=1}^{k}\sum_{l\in\mathcal{I}_{j}}W_{l}^{2}\leq\sum_{l=1}^{n}W_{l}^{2}$.
 Taking the expected value of both sides of (16), it follows that
-$\operatorname{E}[{1\over n}\|X^{n}-{\hat{X}}^{n,{\rm(ML)}}\|_{2}^{2}]$ $\geq q_{0}\eta^{2}\operatorname{E}\left[T\Big{(}1-\sqrt{{1\over T}\sum_{i=1}^{T}W_{i}^{2}}\;\Big{)}^{2}\right]-M^{2}q_{0}\operatorname{E}\left[|{1\over q_{0}}-{n\over k}|(1+{1\over n}\sum_{l=1}^{n}W_{l}^{2})\right]$
-$=q_{0}\eta^{2}\operatorname{E}\left[T\Big{(}1-\sqrt{{1\over T}\sum_{i=1}^{T}W_{i}^{2}}\;\Big{)}^{2}\right]-2M^{2}q_{0}\operatorname{E}\left[|{1\over q_{0}}-{n\over k}|\right],$ (16)
-where the last line follows because the source $\mathbb{X}$ and the speckle noise process are independent. We first focus on the first term. Note that
+$\operatorname{E}[{1\over n}\|X^{n}-{\hat{X}}^{n,{\rm(ML)}}\|_{2}^{2}]$ $\geq q_{0}\eta^{2}\operatorname{E}\left[T\Big{(}1-\sqrt{{1\over T}\sum_{i=1}^{T}W_{i}^{2}}\;\Big{)}^{2}\right]-x_{M}^{2}q_{0}\operatorname{E}\left[|{1\over q_{0}}-{n\over k}|(1+{1\over n}\sum_{l=1}^{n}W_{l}^{2})\right]$
+$=q_{0}\eta^{2}\operatorname{E}\left[T\Big{(}1-\sqrt{{1\over T}\sum_{i=1}^{T}W_{i}^{2}}\;\Big{)}^{2}\right]-2x_{M}^{2}q_{0}\operatorname{E}\left[|{1\over q_{0}}-{n\over k}|\right],$ (17)
+where the last line follows because the source $\mathbf{X}$ and the speckle noise process are independent. We first focus on the first term. Note that
 $\operatorname{E}\left[T\Big{(}1-\sqrt{{1\over T}\sum_{j=1}^{T}w_{j}^{2}}\;\Big{)}^{2}\right]$ $=\operatorname{E}\left[T\Big{(}1+{1\over T}\sum_{j=1}^{T}w_{j}^{2}-2\sqrt{{1\over T_{i}}\sum_{j=1}^{T_{i}}w_{j}^{2}}\Big{)}\right]$
@@
 Therefore, we can bound (18) as
-$2\operatorname{E}\left[T-\sqrt{2T}{\Gamma({T+1\over 2})\over\Gamma({T\over 2})}\right]$ $\geq 2q_{0}(1-{\sqrt{2}\over\Gamma(0.5)})+2q_{0}(1-q_{0})(2-2\Gamma(1.5))+2\operatorname{E}[T(1-{1\over\sqrt{{\rm e}}}{(1-{1\over T})^{T\over 2}\over(1-{2\over T})^{T-1\over 2}})\mathbbmss{1}_{T\geq 3}]$ (21)
+$2\operatorname{E}\left[T-\sqrt{2T}{\Gamma({T+1\over 2})\over\Gamma({T\over 2})}\right]$ $\geq 2q_{0}(1-{\sqrt{2}\over\Gamma(0.5)})+2q_{0}(1-q_{0})(2-2\Gamma(1.5))+2\operatorname{E}[T(1-{1\over\sqrt{{\rm e}}}{(1-{1\over T})^{T\over 2}\over(1-{2\over T})^{T-1\over 2}})\mathds{1}_{T\geq 3}]$ (22)
 Combining (17) and (22), it follows that
 $\operatorname{E}[{1\over n}\|X^{n}-{\hat{X}}^{n,{\rm(ML)}}\|_{2}^{2}]\geq q_{0}$ $\eta^{2}(2q_{0}(1-{\sqrt{2}\over\Gamma(0.5)})+2q_{0}(1-q_{0})(2-2\Gamma(1.5))$ (23)
-$+2\operatorname{E}[T(1-{1\over\sqrt{{\rm e}}}{(1-{1\over T})^{T\over 2}\over(1-{2\over T})^{T-1\over 2}})\mathbbmss{1}_{T\geq 3}])-2M^{2}q_{0}\operatorname{E}\left[|{1\over q_{0}}-{n\over k}|\right].$ (23)
+$+2\operatorname{E}[T(1-{1\over\sqrt{{\rm e}}}{(1-{1\over T})^{T\over 2}\over(1-{2\over T})^{T-1\over 2}})\mathds{1}_{T\geq 3}])-2x_{M}^{2}q_{0}\operatorname{E}\left[|{1\over q_{0}}-{n\over k}|\right].$ (24)
 To bound the last term, note that $k$ is a random variable that depends on $n$. Given $\epsilon>0$, define event $\mathcal{E}$ as
@@
 Using this observation, we can bound the remaining term in (24), as follows:
-$\operatorname{E}\left[|{1\over q_{0}}-{n\over k}|\right]$ $=\operatorname{E}\left[|{1\over q_{0}}-{n\over k}|(\mathbbmss{1}_{\mathcal{E}}+\mathbbmss{1}_{\mathcal{E}^{c}})\right]$
+$\operatorname{E}\left[|{1\over q_{0}}-{n\over k}|\right]$ $=\operatorname{E}\left[|{1\over q_{0}}-{n\over k}|(\mathds{1}_{\mathcal{E}}+\mathds{1}_{\mathcal{E}^{c}})\right]$
 $\leq{\epsilon\over q_{o}(1-\epsilon)}+nP(\mathcal{E}^{c}).$ (26)
 Finally, to bound $P(\mathcal{E}^{c})$, note that
-$P(k(n)\geq nq_{0}(1+\epsilon))$ $=P(\sum_{i=1}^{\lfloor nq_{0}(1+\epsilon)\rfloor}T_{i}<n)$
-$=P\left({1\over\lfloor nq_{0}(1+\epsilon)\rfloor}\sum_{i=1}^{\lfloor nq_{0}(1+\epsilon)\rfloor}T_{i}-{1\over q_{0}}<{n\over\lfloor nq_{0}(1+\epsilon)\rfloor}-{1\over q_{0}}\right)$
-$\leq P\left(|{1\over q_{0}}-{1\over\lfloor nq_{0}(1+\epsilon)\rfloor}\sum_{i=1}^{\lfloor nq_{0}(1+\epsilon)\rfloor}T_{i}|>{1\over q_{0}}-{n\over\lfloor nq_{0}(1+\epsilon)\rfloor}\right)$
+$\operatorname{P}(k(n)\geq nq_{0}(1+\epsilon))$ $=\operatorname{P}(\sum_{i=1}^{\lfloor nq_{0}(1+\epsilon)\rfloor}T_{i}<n)$
+$=\operatorname{P}\left({1\over\lfloor nq_{0}(1+\epsilon)\rfloor}\sum_{i=1}^{\lfloor nq_{0}(1+\epsilon)\rfloor}T_{i}-{1\over q_{0}}<{n\over\lfloor nq_{0}(1+\epsilon)\rfloor}-{1\over q_{0}}\right)$
+$\leq\operatorname{P}\left(|{1\over q_{0}}-{1\over\lfloor nq_{0}(1+\epsilon)\rfloor}\sum_{i=1}^{\lfloor nq_{0}(1+\epsilon)\rfloor}T_{i}|>{1\over q_{0}}-{n\over\lfloor nq_{0}(1+\epsilon)\rfloor}\right)$
 $\leq{\rm e}^{-\lfloor nq_{0}(1+\epsilon)\rfloor\left(q_{0}t_{1}+\ln\frac{1}{1+q_{0}t_{1}}\right)},$ (27)
@@
 Similarly, again using Lemma 2, we have
-$P(k(n)\leq nq_{0}(1-\epsilon))$ $=P(\sum_{i=1}^{\lfloor nq_{0}(1-\epsilon)\rfloor}T_{i}>n)$
-$=P({1\over\lfloor nq_{0}(1-\epsilon)\rfloor}\sum_{i=1}^{\lfloor nq_{0}(1-\epsilon)\rfloor}T_{i}-{1\over q_{0}}>{n\over\lfloor nq_{0}(1-\epsilon)\rfloor}-{1\over q_{0}})$
+$\operatorname{P}(k(n)\leq nq_{0}(1-\epsilon))$ $=\operatorname{P}(\sum_{i=1}^{\lfloor nq_{0}(1-\epsilon)\rfloor}T_{i}>n)$
+$=\operatorname{P}({1\over\lfloor nq_{0}(1-\epsilon)\rfloor}\sum_{i=1}^{\lfloor nq_{0}(1-\epsilon)\rfloor}T_{i}-{1\over q_{0}}>{n\over\lfloor nq_{0}(1-\epsilon)\rfloor}-{1\over q_{0}})$
 $\leq{\rm e}^{-\lfloor nq_{0}(1-\epsilon)\rfloor\left(q_{0}t_{2}+\ln\frac{1}{1+q_{0}t_{2}}\right)},$ (28)
@@
 Combining (26), (27) and (28) with (24) yields the desired result. That is,
-$\operatorname{E}[{1\over n}\|X^{n}-{\hat{X}}^{n}\|_{2}^{2}]\geq$ $q_{0}\eta^{2}(2q_{0}(1-{\sqrt{2}\over\Gamma(0.5)})+2q_{0}(1-q_{0})(2-2\Gamma(1.5))+2\operatorname{E}[T(1-{1\over\sqrt{{\rm e}}}{(1-{1\over T})^{T\over 2}\over(1-{2\over T})^{T-1\over 2}})\mathbbmss{1}_{T\geq 3}])$ (28)
-$-2M^{2}q_{0}\Big{(}{\epsilon\over q_{o}(1-\epsilon)}+n{\rm e}^{-\lfloor nq_{0}(1+\epsilon)\rfloor\left(q_{0}t_{1}+\ln\frac{1}{1+q_{0}t_{1}}\right)}+n{\rm e}^{-\lfloor nq_{0}(1-\epsilon)\rfloor\left(q_{0}t_{2}+\ln\frac{1}{1+q_{0}t_{2}}\right)}\Big{)}.$ (29)
-Proof of Corollary 1.
+$\operatorname{E}[{1\over n}\|X^{n}-{\hat{X}}^{n}\|_{2}^{2}]\geq$ $q_{0}\eta^{2}(2q_{0}(1-{\sqrt{2}\over\Gamma(0.5)})+2q_{0}(1-q_{0})(2-2\Gamma(1.5))+2\operatorname{E}[T(1-{1\over\sqrt{{\rm e}}}{(1-{1\over T})^{T\over 2}\over(1-{2\over T})^{T-1\over 2}})\mathds{1}_{T\geq 3}])$ (29)
+$-2x_{M}^{2}q_{0}\Big{(}{\epsilon\over q_{o}(1-\epsilon)}+n{\rm e}^{-\lfloor nq_{0}(1+\epsilon)\rfloor\left(q_{0}t_{1}+\ln\frac{1}{1+q_{0}t_{1}}\right)}+n{\rm e}^{-\lfloor nq_{0}(1-\epsilon)\rfloor\left(q_{0}t_{2}+\ln\frac{1}{1+q_{0}t_{2}}\right)}\Big{)}.$ (30)
+4.3 Proof of Corollary 1
 For $x\geq 0$ we can show [31] that $\ln(1+x)\leq x\frac{x+6}{2x+6}$ thus,
 $x-\ln(1+x)\geq\frac{x^{2}}{2x+6}.$ (31)
-In addition, we have
+In addition, for $x\geq 0$ we have
 $1-\frac{x}{\lfloor x(1+x^{-1/4})\rfloor}=\Omega(\frac{1}{x^{1/3}}),$ (32)
@@
 Let $\epsilon=(nq_{0})^{-1/4}$ in the result of Theorem 1, using (32), (33) we have $q_{0}t_{1}=\Omega(n^{-1/3})$ and $q_{0}t_{2}=\Omega(n^{-1/3})$, and applying (31) we have
-$\operatorname{E}[{1\over n}\|X^{n}-{\hat{X}}^{n}\|_{2}^{2}]\geq$ $q_{0}\eta^{2}(2q_{0}(1-{\sqrt{2}\over\Gamma(0.5)})+2q_{0}(1-q_{0})(2-2\Gamma(1.5))+2\operatorname{E}[T(1-{1\over\sqrt{{\rm e}}}{(1-{1\over T})^{T\over 2}\over(1-{2\over T})^{T-1\over 2}})\mathbbmss{1}_{T\geq 3}])$
-$-2M^{2}\Big{(}{(nq_{0})^{-1/4}\over 1-(nq_{0})^{-1/4}}+q_{0}n{\rm e}^{-\lfloor nq_{0}(1+(nq_{0})^{-1\over 4})\rfloor\left(c_{1}^{2}n^{-2\over 3}\over 2n^{-1\over 3}+6\right)}+q_{0}n{\rm e}^{-\lfloor nq_{0}(1-(nq_{0})^{-1\over 4})\rfloor\left(c_{2}^{2}n^{-2\over 3}\over 2n^{-1\over 3}+6\right)}\Big{)},$
-where $c_{1},c_{2}\in\mathds{R}^{+}$ are constants. Second line converges to zero as $n\rightarrow\infty$ with $O(n^{-1/4})$. ∎
+$\operatorname{E}[{1\over n}\|X^{n}-{\hat{X}}^{n}\|_{2}^{2}]\geq$ $2q_{0}\eta^{2}\Big{(}q_{0}c_{1}+q_{0}^{2}c_{2}+\operatorname{E}[T(1-{1\over\sqrt{{\rm e}}}{(1-{1\over T})^{T\over 2}\over(1-{2\over T})^{T-1\over 2}})\mathds{1}_{T\geq 3}]\Big{)}$
+$-2x_{M}^{2}\Big{(}{(nq_{0})^{-1/4}\over 1-(nq_{0})^{-1/4}}+q_{0}n{\rm e}^{-\lfloor nq_{0}(1+(nq_{0})^{-1/4})\rfloor\left(d_{1}^{2}n^{-2/3}\over 2n^{-1/3}+6\right)}+q_{0}n{\rm e}^{-\lfloor nq_{0}(1-(nq_{0})^{-1/4})\rfloor\left(d_{2}^{2}n^{-2/3}\over 2n^{-1/3}+6\right)}\Big{)},$
+where $d_{1},d_{2}\in\mathds{R}^{+}$ are constants. Second line converges to zero as $n\rightarrow\infty$ with $O(n^{-1/4})$.
 5 Numerical Experiments
+This section evaluates the performance of the proposed BD-QMAP method.
+We begin by addressing the practical implementation challenges of the BD-QMAP optimization problem.
+Next, we compare its performance against the theoretical lower bound derived in Theorem 1.
+Finally, we investigate BD-QMAP’s application to piecewise-constant sources.
+5.1 Implementation of BD-QMAP
+[…]
+$\operatorname*{arg\,min}_{u^{n}\in\mathcal{X}^{n}_{b}}\frac{1}{n}\sum_{i=k}^{n}\left(\log u_{i}^{2}+\frac{y_{i}^{2}}{u_{i}^{2}}+\frac{\lambda}{b}w_{u_{i-k+1}^{i}}\right),$
+where the term $c_{\mathbf{w}}(u^{n})$ is scaled by $(n-k+1)/n$. This scaling has no impact on the optimal solution, as it can be absorbed into the hyperparameter $\lambda$.
+[…]
+The Viterbi algorithm operates on a trellis diagram with $n$ stages, each containing $|\mathcal{X}_{b}|^{k}$ states.
+A state $s_{i}$ in the diagram corresponds to the subsequence $u_{i-k+1}^{i}$ of length $k$ in the candidate solution.
+The cost associated with state $s_{i}=u^{i}_{i-k+1}$ is updated recursively as:
+$C_{i}(s_{i})=\min_{s^{\prime}\in\mathcal{S}(s_{i})}\left(\log u_{i}^{2}+\frac{y_{i}^{2}}{u_{i}^{2}}+\frac{\lambda}{b}w_{u^{i}_{i-k+1}}+C_{i-1}(s^{\prime})\right),$
+for $i\in\{k,\dots,n\}$, where $\mathcal{S}(s_{i})$ denotes the set of all states that transition into $s_{i}$, and $|\mathcal{S}(s_{i})|=|\mathcal{X}_{b}|$.
+The state $s^{\prime}=u_{i-k}^{i-1}$ represents the preceding state leading to $s_{i}$.
+The BD-QMAPb solution is obtained by minimizing $C_{n}(s_{n})$ recursively.
+5.2 Theoretical Lower Bound
+In Theorem 1, we derived a theoretical lower bound on the minimum achievable reconstruction error for piecewise-constant sources.
+To explore this bound empirically, we compare it with the performance of an ideal despeckler referred to as the genie-aided maximum likelihood (ML) despeckler.
+This ideal estimator assumes access to the exact jump locations in the ground truth signal in addition to the noisy observations.
+With this side information, the reconstruction error is minimized by applying Equation 8 to each constant segment between consecutive jump locations.
+We evaluate this estimator and compare it to the lower bound, along with the experiments on piecewise constant source despeckling using BD-QMAP, as explained in the next section.
+5.3 Comparative analysis of despeckling for piecewise constant sources
+Figure 1: Piecewise constant source ($X$) with parameter $q_{0}=0.001$ sampled under speckle noise ($Y$), enhanced Lee and BD-QMAPb despeckled reconstructions, as discussed in Section 5.
+This subsection evaluates the effectiveness of BD-QMAP in despeckling piecewise-constant sources.
+The results are compared against several well-known despeckling methods, including linear adaptive filters and enhanced versions thereof.
+Additionally, the role of regularization and quantization in BD-QMAP is explored to understand its impact on reconstruction quality and computational efficiency.
+For numerical experiments with the piecewise-constant source introduced in Section 3.2, we assume the absolutely continuous distribution is uniform over the interval $[0,1]$.
+A sample realization of this source under speckle noise is shown in Figure 1.
+For this source, the pivotal parameter affecting despeckler’s performance is $q_{0}$, the probability of a signal level change, which determines the degree of structure in the source [29].
+To train the weights of Equation 2 used in the BD-QMAP regularizer, we sample the random process for $10^{7}$ points with fixed $q_{0}$ and compute the empirical second-order distribution by counting occurrences of $b$-bit quantized sample pairs.
+The BD-QMAP method offers two key practical advantages over existing despecklers.
+First, unlike methods such as [13, 14] and [8], BD-QMAP does not assume an upper bound on speckle power levels in the observations, making it robust to high-noise regimes where other filters fail or bypass the input.
+Second, BD-QMAP is flexible and does not require restrictive assumptions on the signal’s prior distribution, in contrast to Bayesian filters like those in [7, 10, 6].
+These features make BD-QMAP well-suited to a wide range of structured sources.
+Table 1 summarizes the despeckling performance for test datasets of size 100, each containing piecewise-constant signals of length $n=10^{5}$.
+5.3.1 BD-QMAP implementation
 Figure 2: MSE lower bound of the genie-aided ML despeckler as derived in Theorem 1 and the proposed BD-QMAP despecker for piecewise constant source with parameter $q_{0}$.
 ($\operatorname{PSNR}=10\log_{10}(\operatorname{MAX}^{2}/\operatorname{MSE})$ where $\operatorname{MAX}$ is the maximum value of signal’s support set)
-In the following, we first address the practical implementation of the BD-QMAP optimization problem.
-We then compare the performance of BD-QMAP with the lower bound derived in Theorem 1.
-Finally, we perform comparative studies on despeckling for piecewise-constant sources and extending the BD-QMAP to be applied on high dimensional images.
-5.1 Algorithmic considerations
-As noted in Remark 1, addressing the distinct search spaces of the log-likelihood and regularizer in BD-QMAP is essential for practical implementation, and limiting the search to the quantized space ensures feasibility.
-This restriction enables the discrete suboptimal optimization to be efficiently performed using the Viterbi algorithm [32].
-Therefore, to tackle the optimization problem of BD-QMAP, we begin by limiting the search space in (3) to $b$-bit quantized realizations, denoted as $\mathcal{X}^{n}_{b}$.
-We call this suboptimal despeckler BD-QMAPb, as it is optimizing the both terms of (3) in the $b$-bit quantized space.
-In other words, BD-QMAPb solves the following optimization
-$\operatorname*{arg\,min}_{u^{n}\in\mathcal{X}^{n}_{b}}{1\over n}\sum_{i=k}^{n}(\log u_{i}^{2}+\frac{y_{i}^{2}}{u_{i}^{2}}+\frac{\lambda}{b}w_{u_{i-k+1}^{i}}),$ (33)
-where $c_{\mathbf{w}}(u^{n})$ is scaled by $(n-k+1)/n$ without an impact on the optimal solution as it can be absorbed in the hyper-parameter $\lambda$.
-Since similar to the the additive case, the negative log likelihood function of $u^{n}$ is the sum of $n$ element-wise distinct functions, with the $i$-th function only depending on $u_{i}$, the Viterbi algorithm [32] can still be employed to solve (33).
-The Viterbi algorithm operates on a trellis diagram with $n$ stages, each having $|\mathcal{X}_{b}|^{k}$ states.
-State $s_{i}$ in the diagram corresponds to the subsequence $u_{i-k+1}^{i}$ of length $k$ in the candidate solution.
-Let the cost associated to the state $s_{i}=u^{i}_{i-k+1}$ in the diagram follows the recursive rule as,
-$C_{i}(s_{i})=\min_{s^{\prime}\in\mathcal{S}(s_{i})}(\log u_{i}^{2}+\frac{y_{i}^{2}}{u_{i}^{2}}+\frac{\lambda}{b}w_{u^{i}_{i-k+1}})+C_{i-1}(s^{\prime}),$
-for any $i\in\{k,\dots,n\}$, where $\mathcal{S}(s_{i})$ denotes the set of all closest states that transition into the state $s_{i}$ and $|\mathcal{S}(s_{i})|=|\mathcal{X}_{b}|$.
-The state $s^{\prime}=u_{i-k}^{i-1}$ represents the previous state leading to $s_{i}$.
-Thus BD-QMAPb solution can be found recursively starting minimizing $C_{n}(s_{n})$.
-5.2 Theoretical lower bound
-In Theorem 1, a lower bound on the best achievable reconstruction error was derived for the piecewise constant source.
-To explore this bound in practice, we compare it to an ideal despeckler for piecewise constant sources, referred to as the genie-aided maximum likelihood (ML) despeckler.
-This estimator has access to the exact jump locations in the ground truth data in addition to the noisy observations.
-With such side information, the reconstruction error is minimized using Equation 8 for each constant segment between consecutive jump locations.
-For comparison, we evaluated the performance of this genie-aided despeckler against BD-QMAP.
-The BD-QMAP implementation, as detailed in Section 5.3.1, was applied to piecewise constant sources with $q_{0}\in\{0.001,0.01,0.05,0.1\}$.
-Each data point in the figure represents the average error over a test set of 100 realizations with fixed $q_{0}$, each of length $n=10^{6}$.
-5.3 Comparative analysis of despeckling for piecewise constant sources
-Figure 2: Piecewise constant source ($X$) with parameter $q_{0}=0.001$ sampled under speckle noise ($Y$), enhanced Lee and BD-QMAPb despeckled reconstructions, as discussed in Section 5.
-For numerical experiments with the piecewise constant source introduced in Section 3.2 we assume the absolute continous distribution to be uniform on the interval $[0,1]$.
-A sample realization of this source under speckle noise is shown in Figure 2.
-Considering this source, the pivotal parameter which directly affects the performance of the despeckler is $q_{0}$, i.e., the probability of signal level-change which also determines the level of structure for the analog source [29].
-To learn the weights of the Equation 2 to be used in the regularizer term of the BD-QMAP loss function, for a fixed $q_{0}$, we sample the random process for $10^{7}$ points and find the empirical second order distribution by counting the occurances of the $b$ bit quantized samples of length two.
-Key advantages of BD-QMAP over other despecklers in the literature in practice are two fold.
-First, the BD-QMAP assumes no upperbound on the speckle power level on the observation as opposed to [13, 14], and also [8] in which the filter is bypassed completely in high noise level regimes.
-Second, the BD-QMAP method is applicable to any given source without requiring restrictive assumptions on the signal’s prior distribution.
-This flexibility contrasts with the strong assumptions imposed by other Bayesian filters, such as those in [7, 10, 6].
-Each entry in the Table 1 is the despeckling performance for a test dataset of size $100$ of piecewise constant signals of length $n=10^{6}$.
-5.3.1 BD-QMAP implementation
-Using the Viterbi implementation described in Section 5.1 for $k=2$, we can efficiently implement BD-QMAPb.
-As noted in Remark 1, the fidelity term in BD-QMAP optimization is not restricted to a quantized space.
-Therefore, to enhance the BD-QMAPb solution, we refine it by identifying the detected jump locations and computing the ML estimates on the intervals between them using Equation 8.
-This refinement significantly improves performance, especially for highly structured sources (i.e., when $q_{0}$ is small).
-As shown in Table 1, the despeckling performance of BD-QMAPb for highly structured sources improves substantially when the initial estimate is refined for the final BD-QMAP solution.
-Notably, even a low-complexity implementation with $b=2$ is effective in detecting jump locations, with the refinement step yielding significant gains.
+Using the Viterbi implementation described in Section 5.1 for $k=2$, we efficiently implement BD-QMAPb.
+As noted in Remark 1, the fidelity term in BD-QMAP is not restricted to the quantized space.
+Therefore, we refine the BD-QMAPb solution by identifying detected jump locations and applying ML estimation on the intervals between jumps using Equation 8.
+This updated estimation is called BD-QMAP in our experiments.
+This refinement significantly improves performance, particularly for highly structured sources (i.e., small $q_{0}$).
+As shown in Table 1, even a low-complexity implementation with $b=2$ effectively detects jump locations, and the refinement step yields substantial performance gains.
+In Figure 2 the performance of BD-QMAPb and its refined version, BD-QMAP, is compared to the theoretical lower bound from Theorem 1 and the genie-aided ML estimator from Section 5.2.
 5.3.2 Regularization and quantization in BD-QMAP
-To explore the effectiveness of the regularizer of the BD-QMAP in finding the structure of the source, and yeilding lower errors, we conducted experiments for a range of values for $\lambda$ in Equation 3, as shown in Figure 3.
-It can be seen that when $\lambda=0$ the despeckler is not able to distinguish between different levels of structuredness (as measured by $q_{0}$), however a small regularization would let the denoiser to recover the empirical distribution of the ground-truth distribution in the reconstruction.
-Finally, as expected, a large value of $\lambda$ would decrease the fidelity of the reconstruction to the observed noisy signal.
-The number of bits used to quantize the source for the learned weights significantly impacts both the final performance and the time complexity of the optimization problem in Equation 3, as the size of the search space increases exponentially with the number of bits.
-In general, achieving better reconstruction results necessitates selecting the number of bits $b$ in BD-QMAP as high as computationally feasible.
-However, for highly structured data (e.g., high-resolution images), a relatively small number of bits can suffice to uncover the structure of the source from noisy observations.
-As demonstrated in Figure 3, when $q_{0}=0.001$ (indicating a highly structured source), the suboptimal solution of BD-QMAPb with just 2 bits effectively captures the source structure.
-Upon refining the solution to account for quantization error, it achieves a higher PSNR than BD-QMAPb with 3 bits.
-This experiment suggests that for highly structured sources, selecting fewer bits is often sufficient to match the distribution, avoiding the computational challenges associated with higher bit counts.
+To evaluate the impact of the BD-QMAP regularizer, we conduct experiments across a range of values for the hyperparameter $\lambda$ in Equation 3, as shown in Figure 3.
+When $\lambda=0$, the despeckler fails to distinguish between sources with different levels of structure (as measured by $q_{0}$).
+Introducing a small regularization allows the denoiser to recover the empirical distribution of the ground truth in the reconstruction.
+However, excessively large values of $\lambda$ reduce fidelity to the noisy signal.
+The number of bits used for quantization ($b$) also significantly impacts performance and computational complexity.
+While larger $b$ generally yields better reconstructions, for highly structured sources (e.g., high-resolution images), even small $b$ can effectively capture source structure.
+As demonstrated in Figure 3, for $q_{0}=0.001$, the suboptimal BD-QMAPb solution with just $b=2$ bits captures the source structure well.
+After refinement, it achieves higher PSNR than BD-QMAPb with $b=3$, suggesting that smaller $b$ can be sufficient for highly structured data while reducing computational costs.
 Table 1: Despeckling Performance Comparison (PSNR) for Piecewise Constant Source.
@@
 5.3.3 Linear adaptive filters
-All the linear adaptive filters in Table 1 require a window size hyper-parameter that defines the area over which the local statistics will be calculated.
-In these experiments this value is set to $1/2q_{0}$, half of the expected interval lengths of constant pieces in the source.
-We also conduct experiments with an Enhanced version of the filters which is based on the idea of thresholding local variance, as suggested in [8].
-To improve the final performance of the filters, three statiscical classes are considered which are known as heterogeneity adjustments in the literature.
-The simple modification is to distinguish between three region of operation: homogeneous, heterogenous, and strongly speckled.
-To decide in which region the filter operates, if the observation variation coefficient defined as $C_{Y}\triangleq\frac{\operatorname{Var}(Y)}{\operatorname{E}[Y]}$ of the local window falls below a threshold of noise variation coefficient, similarly defined to be $C_{W}\triangleq\frac{\operatorname{Var}(W)}{\operatorname{E}[W]}$, we only do keep the mean window mean value as the despeckled pixel.
-On the other hand, if $C_{Y}\geq C_{\max}$, the filter output is bypassed and the despeckled pixel is equal to the speckled observed pixel, where empirically the optimal value of $C_{\max}$ is set to $\sqrt{3}C_{W}$ [8].
-The Enhanced version of the filters mentioned in Table 1 refer to applying the filter only when $C_{W}\leq C_{Y}\leq C_{\max}$.
+All linear adaptive filters in Table 1 require a window-size hyperparameter to define the region over which local statistics are computed.
+In our experiments, this value is set to $1/2q_{0}$, corresponding to half the expected interval length of constant signal pieces.
+This choice ensures the filters adapt to the level of structure in the source.
+We also experiment with enhanced versions of these filters, incorporating heterogeneity adjustments as proposed in [8].
+This approach classifies regions as homogeneous, heterogeneous, or strongly speckled based on the local coefficient of variation of the observations, defined as $C_{Y}=\operatorname{Var}(Y)/\operatorname{E}[Y]$.
+The enhanced filter operates as follows: If $C_{Y}$ falls below the noise coefficient of variation $C_{W}=\operatorname{Var}(W)/\operatorname{E}[W]$, the filter outputs the mean value of the local window.
+If $C_{Y}\geq C_{\max}$, where $C_{\max}=\sqrt{3}C_{W}$, the filter is bypassed, and the observed pixel is retained.
+Otherwise, the filter is applied normally.
+The enhanced versions of the filters in Table 1 incorporate these modifications to improve despeckling performance.
 6 Conclusion
-A novel Bayesian despeckling method is proposed for structured sources.
-Its performance is analyzed for piecewise constant structured sources, and a lower bound on the minimum achievable MSE is derived.
-The results demonstrate that the despeckler effectively detects edges under speckle noise, achieving superior performance compared to other theoretically grounded approaches.
+A novel Bayesian despeckling method, BD-QMAP, is proposed for structured sources, providing a theoretically grounded approach to addressing speckle noise.
+Its performance is analyzed for piecewise-constant structured sources, with a derived lower bound on the minimum achievable MSE.
+Experimental results highlight the effectiveness of BD-QMAP in despeckling piecewise-constant sources, achieving state-of-the-art performance.
+However, the theoretical understanding of despeckling and its fundamental limits remains significantly less developed compared to denoising in additive noise models, presenting substantial opportunities for further research.
+Additionally, extending the application of QMAP-based methods to image despeckling is a promising direction, which we plan to explore in future work.
 Acknowledgment